A rectangle is 5 feet longer than it is wide, and its area is 66 square feet. Find its dimensions.
Answer by edjones(8007)

L=w+5
L*w=A
w(w+5)=66
w^2+5w=66
w^2+5w-66=0
(w+11)(w-6)=0
w=6 ft
L=6+5=11 ft
Ed